# Symmetry Operators and Gravity

### v1

@@ I Introduction
Quantum field theories often admit a wide class of extended objects, such as solitons, domain walls, defects, and topological operators to name a few. We can study their role in QFT by their explicit insertion in the path integral. Often, this is done at the level of the fields by considering special boundary conditions, or interfaces where the fields undergo non-trivial transformations. This perspective is usually taken in the study of solitonic objects, which are finite energy configurations. From the point of fields, one can study the role of the collective coordinates in the world-volume dynamics of the objects Tong (2005). A particular class of collective coordinates that is always present corresponds to the Goldstone modes associated to the spontaneous symmetry breaking of spacetime symmetries due to the insertion. These are most easily understood from the explicit choices of boundary conditions at the level of the fields. These modes constitute a universal sector of the world-volume dynamics of solitons, and therefore are integral to any effective field theory description of these objects.
@@
The above solitonic picture regulates the insertion of the operator. We first discuss why a choice of regulator is needed in the case of continuous symmetries with Noether currents. Once established, we assume a definition of topological operators in field theory as solitons in general, including those that generate finite symmetries.
@@
In string theory, a quantum theory of gravity, there are studies of symmetries that have shed light on how topological operators can emerge in the decoupling limit of gravity. For example, in García Etxebarria (2022); Apruzzi et al. (2023) it was demonstrated that the topological sector of branes in string theory (i.e., the Wess-Zumino terms in the low energy effective worldvolume action) yield topological operators in field theories constructed both using geometric engineering and AdS/CFT duality. This approach to studying branes as topological operators was further explored in Bah et al. (2024); Apruzzi et al. (2024); Heckman et al. (2022, 2023a); Cvetič et al. (2024); Heckman et al. (2024); Waddleton (2024). The main lesson from these works is that extended non-topological defects in the string theory can become topological operators in the decoupling limit of gravity. In particular, the kinetic terms of the collective coordinates of the branes freeze when $G_{\rm N}$ is taken to zero. This top-down result motivates the present paper in asking if, from a bottom-up perspective, we can see the collective coordinates come alive in a universal way independent from string theory. Our goal is to demonstrate that there is a universal sector, associated to Goldstone modes of spontaneously broken spacetime symmetries, that is necessarily dynamical in presence of gravity.
@@ II Inserting a Symmetry Defect
To illustrate the argument, we begin with a co-dimension one defect $U(\Sigma)$, generating a continuous 0-form symmetry. To keep the notation simple, consider a $U(1)$ symmetry associated to a classically conserved current $d*j=0$. The symmetry action is generated by the topological operator
@@
$\phi(t_{1},\vec{x}_{1})\dots e^{i\alpha\int_{t=0}d^{d-1}x\;j_{0}}\phi(t_{i},\vec{x}_{i})\dots\right>~{},$ (4)
@@
absorbs the defect and transforms the fields with the above phases. After the field redefinition we can take the limit $\lambda\rightarrow 0$ and arrive at the expected result, namely the fact that $U_{\alpha}(\Sigma)$ is topological and can be freely moved around in correlation functions.
@@
$\partial_{\mu}j^{\mu}(x)\Psi(x_{1})\dots\right>=\sum_{i}q_{i}\delta(x-x_{i})\left<\Psi(x_{1})\dots\right>~{},$ (13)
@@
$\partial_{\mu}j^{\mu}(x)\partial^{\prime}_{\nu}j^{\nu}(x^{\prime})\Psi(x_{1})\dots\right>-2\partial^{\prime}_{\mu}\left(\partial^{\mu}\delta(x-x^{\prime})\left<|\phi(x^{\prime})|^{2}\Psi(x_{1})\dots\right>\right)=\sum_{ij}q_{i}q_{j}\delta(x-x_{i})\delta(x^{\prime}-x_{j})\left<\Psi(x_{1})\dots\right>~{}.$
In order to get the expected result, we need to subtract the second term on the left-hand side above, and this is ensured by the presence of the term $|\phi|^{2}$ in (14).
To summarize, to define the defect $U_{\alpha}(\Sigma)$ we gave it a small but non-zero width. This makes the defect into a finite energy-density configuration and thus, in principle, one can study fluctuations of the defect about its classical position. In fact, the defect spontaneously breaks spacetime symmetry and therefore there are Goldstone excitations living on it—these will be the modes of interest when studying its dynamics Low and Manohar (2002).
@@
$\frac{T}{2}\int_{\Sigma}d^{d-1}\sigma\sqrt{\det\frac{\partial x^{\mu}}{\partial\sigma^{a}}\frac{\partial x_{\mu}}{\partial\sigma^{b}}}~{},$ (18)
@@
$\frac{T}{2}\int_{\Sigma}d^{d-1}\sigma(1+(\partial_{a}t_{0})^{2}+\dots)~{},$ (19)
@@
In the zero-width limit $\lambda\rightarrow 0$, the coefficient of the second term approaches a delta function $\delta(t)$; this implies that when integrating over the $t$ direction, we can restrict the integrands to their values at $t=0$, which localizes the resulting action to the worldvolume of the interface. The third term, which on the defect yields a term proportional to $\frac{1}{\lambda}$, is a correction to the bare tension in (19)  [^4] The regulator we are using has the property $f_{\lambda}=\frac{1}{6\lambda}f^{\prime}_{\lambda}-\frac{\lambda}{6}f^{\prime\prime\prime}_{\lambda}$.
@@
where $S_{\text{bulk}}$ is the bulk action in (15), $S_{\rm d}$ is the regulated defect as in (9) and $S_{\rm G}$ is the effective action for the Goldstone. We have also introduced the effective tension
@@ IV Summary and Outlook
In this work we have focused on topological operators for continuous symmetries acting on bosonic fields, realizing them as solitonic objects corresponding to smooth profiles for the fundamental fields. A key feature of continuous symmetries is that they admit Noether currents, which provides a mechanism for obtaining the couplings between the bulk theory and the Goldstones on the defects. We used the interactions between the Goldstone modes and the bulk theory to determine a mechanism by which a defect becomes topological as its width is taken to zero; as we have demonstrated, this automatically imposes the freezing of the Goldstone modes. Moreover, we have used these examples to argue that in the presence of gravity, the zero-width limit of such defects is obstructed due to the emergence of strong gravitational effects that indicate the back-reaction of the defect on the bulk theory. This obstruction offers a complementary perspective on the expected absence of symmetries in gravitational theories. We expect our general conclusions about the existence of defect dynamics to hold in more generic cases (some of which we discuss below), as well as in cases for which gravitational interactions are present from the beginning of the analysis.
@@
Recent work Apruzzi et al. (2023); García Etxebarria (2022); Heckman et al. (2023b, 2024) has explored the existence of topological operators in holographic theories. These cases have been a strong motivation for analyzing the fate of topological operators in QFTs coupled to gravity in terms of the dynamics of soliton-like objects. In these works, it was demonstrated in the context of AdS/CFT that such objects can become topological on the conformal boundary of AdS, leading to the proposal that branes “are” symmetry operators. This is consistent with our discussion since $\kappa$ is renormalized to zero on the boundary. In light of our perturbative field theory analysis, we also expect that there exists a regulated version of an extended dynamical object in AdS that, when pushed to the conformal boundary, becomes a regulated version of a topological operator in the holographic dual field theory. Thus, another natural next step is to carefully work out a suitable version of a regulated worldvolume action for branes in the AdS bulk that is compatible with our field theory analysis when pushed to the conformal boundary.
@@ Acknowledgments
We are grateful to Zohar Komargodski for comments on the manuscript. We would like to thank Fabio Apruzzi, Nima Arkani-Hamed, Federico Bonetti, Thomas Dumitrescu, Jonathan Heckman, Simeon Hellerman, Ken Intriligator, David E. Kaplan, Shlomo Razamat, Sakura Schafer-Nameki, Yuji Tachikawa, Cumrun Vafa, and Amos Yarom for illuminating discussions. PJ is supported by the Johns Hopkins University Provost’s Postdoctoral Fellowship. IB, KR, PJ, and TW are supported in part by the Simons Collaboration on Global Categorical Symmetries and also by the NSF grant PHY-2112699.

### v2

@@ I Introduction
Quantum field theories often admit a wide class of extended objects, such as solitons, domain walls, defects, and topological operators to name a few. We can study their role in QFT by their explicit insertion in the path integral. Often, this is done at the level of the fields by considering special boundary conditions, or interfaces where the fields undergo non-trivial transformations. This perspective is usually taken in the study of solitonic objects, which are finite energy configurations. From the point of view of fields, one can study the role of the collective coordinates in the world-volume dynamics of such solitonic objects Tong (2005). A particular class of collective coordinates that is always present corresponds to the Goldstone modes associated to the spontaneous breaking of spacetime symmetries due to the insertion. These are most easily understood in terms of the explicit choices of boundary conditions imposed on the fields. These modes constitute a universal sector of the world-volume dynamics of solitons, and therefore are integral to any effective field theory description of these objects.
@@
In string theory, a quantum theory of gravity, there are studies of symmetries that have shed light on how topological operators can emerge in the decoupling limit of gravity. For example, in Garc´ıa Etxebarria (2022); Apruzzi et al. (2023) it was demonstrated that the topological sector of branes in string theory (i.e., the Wess-Zumino terms in the low energy effective worldvolume action) yield topological operators in field theories constructed both using geometric engineering and AdS/CFT duality. This approach to studying branes as topological operators was further explored in Bah et al. (2024); Apruzzi et al. (2024); Heckman et al. (2022, 2023a); Cvetič et al. (2024); Heckman et al. (2024); Waddleton (2024). The main lesson from these works is that extended non-topological defects in the string theory can become topological operators in the decoupling limit of gravity. In particular, the kinetic terms of the collective coordinates of the branes freeze when $G_{\rm N}$ is taken to zero.
This top down perspective inspires our main approach in this paper, as it suggests that we can analyze topological operators when coupling field theories to gravity by treating them as solitons. While it would be interesting to provide a purely field-theoretic justification for this assumption, we settle with this motivation from string theory. With this assumption, our aim is to determine whether or not we can see the collective coordinates come alive in a universal fashion when gravity is turned on. Said differently, our goal is to demonstrate that there is a universal sector, associated to Goldstone modes of spontaneously broken spacetime symmetries, that is necessarily dynamical in presence of gravity.
It is also worth pointing out that in a theory of gravity, the geometry (and presumably also the topology) of the spacetime background fluctuates. We therefore expect the dynamics of the collective coordinates to be induced by these fluctuations, and furthermore that the dynamics of the collective coordinates freeze or decouple in the field theory limit.
@@
The above solitonic picture regulates the insertion of the operator. We first discuss why a choice of regulator is needed in the case of continuous symmetries with Noether currents (see Cherman et al. (2022) for a related discussion). Once this point is established, we assume a definition of topological operators in field theory as solitons in general, including those that generate finite symmetries.
@@ II Inserting a Symmetry Defect
To illustrate the argument, we begin with a co-dimension one defect $U(\Sigma)$, generating a continuous 0-form symmetry. To keep the notation simple, consider a $U(1)$ symmetry associated to a classically conserved current dj=0d*j=0. The symmetry action is generated by the topological operator
@@
$\left<\phi(t_{1},\vec{x}_{1})\cdots e^{i\alpha\int_{t=0}d^{d-1}x\;j_{0}}\phi(t_{i},\vec{x}_{i})\cdots\right$>~, (4)
@@
absorbs the defect and transforms the fields with the above phases  [^4] Note that the Jacobian determinant associated to the spacetime-dependent field redefinition of the complex scalar in (10) is trivial.. After the field redefinition, we can take the limit $\lambda\rightarrow 0$ and arrive at the expected result, namely the fact that $U_{\alpha}(\Sigma)$ is topological and can be freely moved around in correlation functions.
@@
$\left<\partial_{\mu}j^{\mu}(x)\Psi(x_{1})\cdots\right>=\sum_{i}q_{i}\delta(x-x_{i})\left<\Psi(x_{1})\cdots\right$>~, (13)
@@
$\left<\partial_{\mu}j^{\mu}(x)\partial^{\prime}_{\nu}j^{\nu}(x^{\prime})\Psi(x_{1})\cdots\right>-2\partial^{\prime}_{\mu}\left(\partial^{\mu}\delta(x-x^{\prime})\left<|\phi(x^{\prime})|^{2}\Psi(x_{1})\cdots\right>\right)=\sum_{ij}q_{i}q_{j}\delta(x-x_{i})\delta(x^{\prime}-x_{j})\left<\Psi(x_{1})\cdots\right$>~.
In order to get the expected result, namely the charge of field-insertions after integrating the divergences of the currents as in (14), we need to subtract the second term on the left-hand side above, and this is ensured by the presence of the term $|\phi|^{2}$ in (14).
To summarize, to define the defect $U_{\alpha}(\Sigma)$ we gave it a small but non-zero width. This causes the field to transform smoothly across the defect. Motivated by this picture as well as the top down perspective in holography, we are going to view a symmetry defect as a solitonic object which can fluctuate around its classical position. The insertion of this solitonic object spontaneously breaks spacetime symmetry, and the fluctuations correspond to the dynamics of the resulting Goldstone excitations living on the soliton—these will be the modes of interest when studying its dynamics Low and Manohar (2002).
@@
$T\int_{\Sigma}d^{d-1}\sigma\sqrt{\det\frac{\partial x^{\mu}}{\partial\sigma^{a}}\frac{\partial x_{\mu}}{\partial\sigma^{b}}}~,$ (18)
@@
$\frac{T}{2}\int_{\Sigma}d^{d-1}\sigma(1+(\partial_{a}t_{0})^{2}+\cdots)~,$ (19)
@@
In the zero-width limit $\lambda\rightarrow 0$, the coefficient of the second term approaches a delta function $\delta(t)$; this implies that when integrating over the $t$ direction, we can restrict the integrands to their values at $t=0$, which localizes the resulting action to the worldvolume of the interface. The third term, which on the defect yields a term proportional to $\frac{1}{\lambda}$, is a correction to the bare tension in (19)  [^5] The regulator we are using has the property $f_{\lambda}^{\prime 2}=\frac{1}{6\lambda}f^{\prime}_{\lambda}-\frac{\lambda}{6}f^{\prime\prime\prime}_{\lambda}$.
@@
where $S_{\text{bulk}}$ is the bulk action in (15), $S_{\rm d}$ is the regulated defect as in (9) and $S_{\rm G}$ is the effective action for the Goldstone mode. We have also introduced the effective tension
@@ IV Summary and Outlook
In this work we have focused on topological operators for continuous symmetries acting on bosonic fields, realizing them as solitonic objects corresponding to smooth profiles for the fundamental fields. A key feature of continuous symmetries is that they admit Noether currents, which provides a mechanism for obtaining the couplings between the bulk theory and the fluctuations of the collective coordinates on the defects. We used the interactions between the collective coordinates and the bulk theory to determine a mechanism by which a defect becomes topological as its width is taken to zero; as we have demonstrated, this automatically imposes the freezing of the associated Goldstone modes. Moreover, we have used these examples to argue that in the presence of gravity, the zero-width limit of such defects is obstructed due to the emergence of strong gravitational effects that indicate the back-reaction of the defect on the bulk theory. This obstruction offers a complementary perspective on the expected absence of symmetries in gravitational theories. We expect that our general conclusions about the existence of defect dynamics to hold in more generic cases (some of which we discuss below), as well as in cases for which gravitational interactions are present from the beginning of the analysis.
@@
Recent work Apruzzi et al. (2023); Garc´ıa Etxebarria (2022); Heckman et al. (2023b, 2024) has explored the existence of topological operators in holographic theories. These cases have been a strong motivation for analyzing the fate of topological operators in QFTs coupled to gravity in terms of the dynamics of soliton-like objects. In these works, it was demonstrated in the context of AdS/CFT that such objects can become topological on the conformal boundary of AdS, leading to the proposal that branes “are” symmetry operators. This is consistent with our discussion since $\kappa$ is renormalized to zero on the boundary. In light of our perturbative field theory analysis, we also expect that there exists a regulated version of an extended dynamical object in AdS that, when pushed to the conformal boundary, becomes a regulated version of a topological operator in the holographic dual field theory. Thus, another natural next step is to carefully work out a suitable version of a regulated worldvolume action for branes in the AdS bulk that is compatible with our field theory analysis when pushed to the conformal boundary.
@@ Acknowledgments
We are grateful to Zohar Komargodski for comments on the manuscript. We would like to thank Fabio Apruzzi, Nima Arkani-Hamed, Federico Bonetti, Thomas Dumitrescu, Jonathan Heckman, Simeon Hellerman, Ken Intriligator, David E. Kaplan, Shlomo Razamat, Sakura Schafer-Nameki, Yuji Tachikawa, Cumrun Vafa, and Amos Yarom for illuminating discussions. PJ is supported by the Johns Hopkins University Provost’s Postdoctoral Fellowship. IB, KR, PJ, and TW are supported in part by the Simons Collaboration on Global Categorical Symmetries and also by the NSF grant PHY-2412361.